OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 5. Find y when x = 10.

OpenStudy (anonymous):

y = kx^2 5 = k(5)^2 5 = 25k k = 1/5 k is the constant by the way. y = (1/5)(10)^2 y = 20